OpenStudy (anonymous):

Use Euler's Formula to find the missing number. Faces:_____ Edges:15 Vertices: 9

OpenStudy (decentnabeel):

do you know Euler's Formula @NiraQ

OpenStudy (decentnabeel):

the Euler's formula is.. V - E + F = 2 V is the vertices E is the edges F is the faces V - E + F = 2 subtract V and add E on both sides F=2-V+E plug the value F=2-9+15 F=8 the answer is.. faces =8

OpenStudy (decentnabeel):

Are you understand @NiraQ